# A quantum machine learning classifier to search for new physics

### v4

@@ 3.4.3 Numerical results
After the $N_{part}$ cut, the ansatz is trained. The fidelities for $l=5$, $10$ and $15$ ansatz are $12\%, 47.47$$47\%$ and $47.53\%$, respectively. A validation set is made which is consist of events after $N_{part}$ cut with $73400$ and $108000$ NP and SM events (all different from the training set), respectively. The normalized distributions of $\hat{a}$ and $\hat{b}$ for the validation set are shown in Fig. 13. It can be seen that, for both VQSN and t-KNN event selection strategies, $\hat{a}$ and $\hat{b}$ can be used for discriminate the signal events from background events. The distributions of $\hat{a}$ for the VQSN shows that the difference from different $l$s is small. Furthermore, the $\hat{a}$ of a background event is more closely aligned with 0 when using an ansatz.
@@
where the summation $\sum_{p}$ is over the three processes $p\to jj\ell^{+}\ell^{-}\nu\bar{\nu}, pp$$p\to j\ell^{+}\ell^{-}\nu\bar{\nu}$ and $pp\to\gamma j\ell^{+}\ell^{-}\nu\bar{\nu}$, $\sigma_{t\bar{t}}$ is the inclusive cross-section of the production of $t\bar{t}+jets$ which is taken as $832\;({\rm pb})$, $\varepsilon_{b}=1-77\%$ is from mis-tagging of $b$-jets, $Br_{W\to\ell\nu}=21.7\%$ ParticleDataGroup:2020ssz , $\varepsilon^{t\bar{t}}_{N_{part}}$ is the cut efficiency of $N_{part}$ cut in Table 6, $\varepsilon_{t\bar{t}}$ is the cut efficiency of (V)QSN and (t-)KNN estimated using the $t\bar{t}$ events, and $\tilde{\sigma}_{bg}$ is the sum of $\ell\ell+jets$, $j\ell\ell\ell\nu$, and $\gamma\ell\ell+jets$ backgrounds after all cuts. Then, $\sigma_{bg}=\sigma(f_{0}=0)$ and $\sigma_{s}=\sigma(f_{0})-\sigma_{bg}$,
@@
Figure 15: Comparison of the constraints on $M_{0}$ when $f_{0}$ is positive at $\mathcal{S}_{stat}=2$ obtained from different event selection strategies. The result of Ref. Ellis:2018cos is shown as the dashed line.
Using the upper bound of coefficient constraint, one can set a constraint on $M_{0}$ when $f_{0}$ is positive. The results at $\mathcal{S}_{stat}=2$ are shown in Fig. 15, which provides a more clear comparison of different algorithms. It can be seen that t-KNN ($k=1000$) is the most effective, followed by t-KNN ($k=100$), VQSN ($l=15$), and VQSN ($l=10$). Furthermore, the result presented in Ref. Ellis:2018cos  ($M_{0}\geq 1040\;{\rm MeV}$) are shown as a dashed line, and the four aforementioned cases demonstrate tighter constraints. This suggests that the processes under consideration are sensitive to the gQGCs and are worth of consideration in the global fit of the constraints on gQGCs. Furthermore, the results demonstrate the efficacy of the t-KNN and VQSN algorithms employed. The validity is a crucial consideration when using an EFT Dong:2023nir ; Yang:2021pcf ; Guo:2020lim ; Yue:2021snv ; Fu:2021mub ; Yang:2022ilt ; Layssac:1993vfp ; Corbett:2017qgl ; Perez:2018kav ; Almeida:2020ylr ; Kilian:2018bhs ; Kilian:2021whd , and can usually be tested through unitarity. According to the Ref. Ellis:2018cos , the region of coefficients in this study is free from the violation of unitarity.
@@
3.5 Search for the anomalies quartic gauge couplings in the tri-photon process at muon colliders
3.5.1 The contribution of the anomalies quartic gauge couplings in the tri-photon process
As machine learning algorithms, (V)QSN and (t-)KNN should not depend on the NP or process to be searched. Thus, as an example, we consider the dimension-8 operators contributing to the anomalies quartic gauge couplings (aQGCs) in the tri-photon process at muon colliders. Due to the attention paid to vector boson scattering (VBS) processes at the LHC, the aQGCs which are suitable to be studied in VBS processes have also been widely studied phenomenologically Green:2016trm ; Chang:2013aya ; Anders:2018oin ; Zhang:2018shp ; Bi:2019phv and experimentally ATLAS:2014jzl ; CMS:2020gfh ; ATLAS:2017vqm ; CMS:2017rin ; CMS:2020ioi ; CMS:2016gct ; CMS:2017zmo ; CMS:2018ccg ; ATLAS:2018mxa ; CMS:2019uys ; CMS:2016rtz ; CMS:2017fhs ; CMS:2019qfk ; CMS:2020ypo ; CMS:2020fqz .
@@ 3.5.1 The contribution of the anomalies quartic gauge couplings in the tri-photon process
The dimension-8 operators contributing to the aQGCs are often classified as scalar/longitudinal operators $O_{S_{i}}$, mixed transverse and longitudinal operators $O_{M_{i}}$ and transverse operators $O_{T_{i}}$. It has been found that the tri-photon processes at the muon colliders are sensitive to the $O_{T_{i}}$ operators Yang:2020rjt , they are Eboli:2006wa ; Eboli:2016kko ; Anders:2018oin ,
@@
where $\widehat{W}\equiv\vec{\sigma}\cdot{\vec{W}}/2$ with $\sigma$ being the Pauli matrices and ${\vec{W}}=\{W^{1},W^{2},W^{3}\}$, $B_{\mu}$ and $W_{\mu}^{i}$ are $U(1)_{\rm Y}$ and $SU(2)_{\rm I}$ gauge fields, and $B_{\mu\nu}$ and $W_{\mu\nu}$ correspond to the field strength tensors. The Feynman diagrams are shown in Fig. 16. The contribution of $O_{T,1}$ and $O_{T,6}$ operator are as same as the ones of $O_{T,0}$ and $O_{T,5}$, respectively, therefore are not studied.
@@ 3.5.2 Training of the ansatz
Different from the previous subsections, in the generation of the events, the standard cuts are set to be as same as the default except for the $p_{\gamma}^{T}$ cut which is set to be $p_{\gamma}^{T}\geq 0.1E_{beam}$ where $E_{beam}$ is the beam energy which can suppressing the SM background and ease the pressure on numerical computation consumption. The cases of $\sqrt{s}=3\;{\rm TeV}$, $10\;{\rm TeV}$, $14\;{\rm TeV}$ and $30\;{\rm TeV}$ are considered Black:2022cth ; Accettura:2023ked . To generate the training set, $2^{16}+2^{16}$ events are chosen from the SM background and the NP events. It has been shown that, the kinematic features are similar for different operators, therefore in the training set, we use only the events from $O_{T_{0}}$. It is required that the final state to have at least three photons. The observables chosen to form the feature space are listed in Table 10, where $E_{\gamma_{1,2,3}}$ and $p^{T}_{\gamma_{1,2,3}}$ are energies and transverse momenta of the hardest, second hardest and the third hardest photons. Using this feature space, $d=6$ and $d^{\prime}=4$, and with $2^{17}$ events, the ansatz takes $n_{q}=19$ qubits. In this subsection, we also use the SCA entanglement layer and with $l=10$, the circuit is the one shown in Fig. 4. After training, the fidelities are $\%, 46.846.8\%$, $46.9\%$ and $46.9\%$ for the cases of $\sqrt{s}=3\;{\rm TeV}$, $10\;{\rm TeV}$, $14\;{\rm TeV}$ and $30\;{\rm TeV}$, respectively.
@@ 3.5.3 Numerical results
Using the designed luminosities for both the “conservative” and “optimistic” cases Black:2022cth ; Accettura:2023ked and the fitted cross-sections, the expected constraints can be estimated using Eq. (4). The results are listed in Tables 12 and 13. It can be shown that the VQSN is effective in the search for aQGCs, the expected constraints generally outperform the ones obtained by a traditional event selection strategy Yang:2020rjt .
@@ 3.6 Discussion on efficacy difference between (t-)KNN and (V)QSN
While the VQSN method can technically be adapted to classify using only nearest neighbors basheer2021quantumknearestneighborsalgorithm , implementing this quantum variant demands additional gate operations. This not only fails to surpass the computational efficiency of classical t-KNN when processing classical data but also produces deeper circuits impractical for execution on current noisy quantum hardware. Consequently, we contend that a quantum algorithm achieving coefficient constraints on par with t-KNN (within the same order of magnitude) while simultaneously reducing cc by an order of magnitude and remaining executable on current quantum hardware would hold significant value for high-luminosity collider experiments confronting exponentially growing datasets.
@@ 4 Summary
The VQSN is used in the phenomenological study of the gQGCs in the process $\mu^{+}\mu^{-}\to jj\nu\bar{\nu}$ at a $\sqrt{s}=3$ TeV muon collider. The VQSN event selection strategy is tested on the tianyan176-2. The expected coefficient constraints are obtained and found to be at the same order of magnitude as the traditional event selection strategy in Ref. Yang:2023gos , indicating that the VQSN is an effective method and possesses noise tolerance.
The gQGCs at the LHC are also studied, using the processes $p\to jj\ell^{+}\ell^{-}\nu\bar{\nu}, pp$$p\to j\ell^{+}\ell^{-}\nu\bar{\nu}$ and $pp\to\gamma j\ell^{+}\ell^{-}\nu\bar{\nu}$. These processes can be concluded to be sensitive to the gQGCs. The constraints on the coefficient are of a similar magnitude to those obtained in the $pp\to\gamma\gamma$ process Ellis:2018cos . Consequently, they can serve as a valuable supplement to the constraints on the gQGCs. Moreover, the $pp\to jj\ell^{+}\ell^{-}\nu\bar{\nu}$ process contains a VBS process which is expected to be more significant at higher energies Yang:2020rjt , therefore is expected to be more important at future colliders such as the HE-LHC and FCC-hh FCC:2018bvk ; Benedikt:2022kan .
Another application is the contribution of the aQGCs in the tri-photon processes at muon colliders. In this case, only the VQSN is studied. The expected constraints are obtained, and found to be more strengthen than the ones obtained using an traditional event selection strategy Yang:2020rjt .
@@
This study has not addressed further improvements in encoding the datasets in optimistic ways Havlicek:2018nqz ; Lloyd:2020eeh which are suitable for cooperation with (V)QSN. Furthermore, it would be beneficial to investigate the simulation of the collision process on a quantum computer. Reading results directly from a quantum computer is resource intensive, QSN can process data within the quantum computer, avoiding this consumption. In particular, it is plausible that the training dataset originates from simulations on a quantum computer, while the test dataset is derived from classical data obtained from a collider. This scenario is well-suited to the QSN.

### v5

@@ 3.4.3 Numerical results
After the $N_{part}$ cut, the ansatz is trained. The fidelities for $l=5$, $10$ and $15$ ansatz are $12\%, 47.47$$47\%$ and $47.53\%$, respectively. A validation set is made which consists of events after $N_{part}$ cut with $73400$ and $108000$ NP and SM events (all different from the training set), respectively. The normalized distributions of $\hat{a}$ and $\hat{b}$ for the validation set are shown in Fig. 13. It can be seen that, for both VQSN and t-KNN event selection strategies, $\hat{a}$ and $\hat{b}$ can be used for discriminating the signal events from background events. The distributions of $\hat{a}$ for the VQSN show that the difference from different $l$s is small. Furthermore, the $\hat{a}$ of a background event is more closely aligned with 0 when using an ansatz.
@@
where the summation $\sum_{p}$ is over the three processes $p\to jj\ell^{+}\ell^{-}\nu\bar{\nu}, pp$$p\to j\ell^{+}\ell^{-}\nu\bar{\nu}$ and $pp\to\gamma j\ell^{+}\ell^{-}\nu\bar{\nu}$, $\sigma_{t\bar{t}}$ is the inclusive cross-section of the production of $t\bar{t}+jets$ which is taken as $832\;({\rm pb})$, $\varepsilon_{b}=1-77\%$ is from mis-tagging of $b$-jets, $Br_{W\to\ell\nu}=21.7\%$ Zyla and others (2020), $\varepsilon^{t\bar{t}}_{N_{part}}$ is the cut efficiency of $N_{part}$ cut in Table 6, $\varepsilon_{t\bar{t}}$ is the cut efficiency of (V)QSN and (t-)KNN estimated using the $t\bar{t}$ events, and $\tilde{\sigma}_{bg}$ is the sum of $\ell\ell+jets$, $j\ell\ell\ell\nu$, and $\gamma\ell\ell+jets$ backgrounds after all cuts. Then, $\sigma_{bg}=\sigma(f_{0}=0)$ and $\sigma_{s}=\sigma(f_{0})-\sigma_{bg}$,
@@
Figure 15: Comparison of the constraints on $M_{0}$ when $f_{0}$ is positive at $\mathcal{S}_{stat}=2$ obtained from different event selection strategies. The result of Ref. Ellis and Ge (2018) is shown as the dashed line.
Using the upper bound of coefficient constraint, one can set a constraint on $M_{0}$ when $f_{0}$ is positive. The results at $\mathcal{S}_{stat}=2$ are shown in Fig. 15, which provides a more clear comparison of different algorithms. It can be seen that t-KNN ($k=1000$) is the most effective, followed by t-KNN ($k=100$), VQSN ($l=15$), and VQSN ($l=10$). Furthermore, the result presented in Ref. Ellis and Ge (2018) ($M_{0}\geq 1040\;{\rm MeV}$) is shown as a dashed line, and the four aforementioned cases demonstrate tighter constraints. This suggests that the processes under consideration are sensitive to the gQGCs and are worthy of consideration in the global fit of the constraints on gQGCs. Furthermore, the results demonstrate the efficacy of the t-KNN and VQSN algorithms employed. The validity is a crucial consideration when using an EFT Dong et al. (2023); Yang et al. (2021); Guo et al. (2020a); Yue et al. (2023); Fu et al. (2021); Yang et al. (2023b); Layssac et al. (1994); Corbett et al. (2017); Perez et al. (2018); Almeida et al. (2020); Kilian et al. (2020, 2021), and can usually be tested through unitarity. According to the Ref. Ellis and Ge (2018), the region of coefficients in this study is free from the violation of unitarity.
@@
3.5 Search for the anomalous quartic gauge couplings in the tri-photon process at muon colliders
3.5.1 The contribution of the anomalous quartic gauge couplings in the tri-photon process
As machine learning algorithms, (V)QSN and (t-)KNN should not depend on the NP or process to be searched. Thus, as an example, we consider the dimension-8 operators contributing to the anomalous quartic gauge couplings (aQGCs) in the tri-photon process at muon colliders. Due to the attention paid to vector boson scattering (VBS) processes at the LHC, the aQGCs which are suitable to be studied in VBS processes have also been widely studied phenomenologically Green et al. (2017); Chang et al. (2013); Anders and others (2018); Zhang and Zhou (2019); Bi et al. (2019) and experimentally Aad and others (2014); Sirunyan and others (2020c); Aaboud and others (2017); Khachatryan and others (2017b); Sirunyan and others (2020a); Khachatryan and others (2017a); Sirunyan and others (2017, 2019c); Aaboud and others (2019); Sirunyan and others (2019a); Khachatryan and others (2016); Sirunyan and others (2018, 2019b, 2020d, 2021).
@@ 3.5.1 The contribution of the anomalous quartic gauge couplings in the tri-photon process
The dimension-8 operators contributing to the aQGCs are often classified as scalar/longitudinal operators $O_{S_{i}}$, mixed transverse and longitudinal operators $O_{M_{i}}$ and transverse operators $O_{T_{i}}$. It has been found that the tri-photon processes at the muon colliders are sensitive to the $O_{T_{i}}$ operators Yang et al. (2020), they are Eboli et al. (2006); Éboli and Gonzalez-Garcia (2016); Anders and others (2018),
@@
where $\widehat{W}\equiv\vec{\sigma}\cdot{\vec{W}}/2$ with $\sigma$ being the Pauli matrices and ${\vec{W}}=\{W^{1},W^{2},W^{3}\}$, $B_{\mu}$ and $W_{\mu}^{i}$ are $U(1)_{\rm Y}$ and $SU(2)_{\rm I}$ gauge fields, and $B_{\mu\nu}$ and $W_{\mu\nu}$ correspond to the field strength tensors. The Feynman diagrams are shown in Fig. 16. The contributions of $O_{T,1}$ and $O_{T,6}$ operator are as same as the ones of $O_{T,0}$ and $O_{T,5}$, respectively, therefore are not studied.
@@ 3.5.2 Training of the ansatz
Different from the previous subsections, in the generation of the events, the standard cuts are set to be as same as the default except for the $p_{\gamma}^{T}$ cut which is set to be $p_{\gamma}^{T}\geq 0.1E_{beam}$ where $E_{beam}$ is the beam energy which can suppress the SM background and ease the pressure on numerical computation consumption. The cases of $\sqrt{s}=3\;{\rm TeV}$, $10\;{\rm TeV}$, $14\;{\rm TeV}$ and $30\;{\rm TeV}$ are considered Black and others (2024); Accettura and others (2023). To generate the training set, $2^{16}+2^{16}$ events are chosen from the SM background and the NP events. It has been shown that, the kinematic features are similar for different operators, therefore in the training set, we use only the events from $O_{T_{0}}$. It is required that the final state to have at least three photons. The observables chosen to form the feature space are listed in Table 10, where $E_{\gamma_{1,2,3}}$ and $p^{T}_{\gamma_{1,2,3}}$ are energies and transverse momenta of the hardest, second hardest and the third hardest photons. Using this feature space, $d=6$ and $d^{\prime}=4$, and with $2^{17}$ events, the ansatz takes $n_{q}=19$ qubits. In this subsection, we also use the SCA entanglement layer and with $l=10$, the circuit is the one shown in Fig. 4. After training, the fidelities are $\%, 46.846.8\%$, $46.9\%$ and $46.9\%$ for the cases of $\sqrt{s}=3\;{\rm TeV}$, $10\;{\rm TeV}$, $14\;{\rm TeV}$ and $30\;{\rm TeV}$, respectively.
@@ 3.5.3 Numerical results
Using the designed luminosities for both the “conservative” and “optimistic” cases Black and others (2024); Accettura and others (2023) and the fitted cross-sections, the expected constraints can be estimated using Eq. (4). The results are listed in Tables 12 and 13. It can be shown that the VQSN is effective in the search for aQGCs, the expected constraints generally outperform the ones obtained by a traditional event selection strategy Yang et al. (2020).
@@ 3.6 Discussion on efficacy difference between (t-)KNN and (V)QSN
While the VQSN method can technically be adapted to classify using only nearest neighbors Basheer et al. (2021), implementing this quantum variant demands additional gate operations. This not only fails to surpass the computational efficiency of classical t-KNN when processing classical data but also produces deeper circuits impractical for execution on current noisy quantum hardware. Consequently, we contend that a quantum algorithm achieving coefficient constraints on par with t-KNN (within the same order of magnitude) while simultaneously reducing cc by an order of magnitude and remaining executable on current quantum hardware would hold significant value for high-luminosity collider experiments confronting exponentially growing datasets.
@@ 4 Summary
The VQSN is used in the phenomenological study of the gQGCs in the process $\mu^{+}\mu^{-}\to jj\nu\bar{\nu}$ at a $\sqrt{s}=3$ TeV muon collider. The VQSN event selection strategy is tested on the tianyan176-2. The expected coefficient constraints are obtained and found to be at the same order of magnitude as the traditional event selection strategy in Ref. Yang et al. (2023a), indicating that the VQSN is an effective method and possesses noise tolerance.
The gQGCs at the LHC are also studied, using the processes $p\to jj\ell^{+}\ell^{-}\nu\bar{\nu}, pp$$p\to j\ell^{+}\ell^{-}\nu\bar{\nu}$ and $pp\to\gamma j\ell^{+}\ell^{-}\nu\bar{\nu}$. These processes can be concluded to be sensitive to the gQGCs. The constraints on the coefficient are of a similar magnitude to those obtained in the $pp\to\gamma\gamma$ process Ellis and Ge (2018). Consequently, they can serve as a valuable supplement to the constraints on the gQGCs. Moreover, the $pp\to jj\ell^{+}\ell^{-}\nu\bar{\nu}$ process contains a VBS process which is expected to be more significant at higher energies Yang et al. (2020), therefore is expected to be more important at future colliders such as the HE-LHC and FCC-hh Abada and others (2019); Benedikt and others (2022).
Another application is the contribution of the aQGCs in the tri-photon processes at muon colliders. In this case, only the VQSN is studied. The expected constraints are obtained, and found to be more strengthened than the ones obtained using a traditional event selection strategy Yang et al. (2020).
@@
This study has not addressed further improvements in encoding the datasets in optimistic ways Havlicek et al. (2019); Lloyd et al. (2020) which are suitable for cooperation with (V)QSN. Furthermore, it would be beneficial to investigate the simulation of the collision process on a quantum computer. Reading results directly from a quantum computer is resource intensive, QSN can process data within the quantum computer, avoiding this consumption. In particular, it is plausible that the training dataset originates from simulations on a quantum computer, while the test dataset is derived from classical data obtained from a collider. This scenario is well-suited to the QSN.